If f(z)= 1/\sqrt{|z-1} find the domain of definition of f.


1
Expert's answer
2021-08-20T04:51:56-0400
"f(z)=\frac{1}{\sqrt{z-1}}"

Domain for function


"f(x)=\frac{g(x)}{h(x)}" is equal to "h(x)\neq 0."

Thus we will have


"\sqrt{z-1}\neq 0,\newline
z-1\neq 0,\newline
z\neq 1."

Answer. "(-\infty,1)\bigcup (1,\infty)."